For the graph on the trade action row the options are for Corn Dolorium exports 26 or imports 26. For jeans Dolorium the options are exports 78 or imports 78. For corn Arcadia the options are exports 26 or imports 26. For jeans Arcadia the options are exports 78 or imports 78.
3. Gains from trade Consider two neighboring island countries called Dolorium and Arcadia. They each have 4 miltion labor hours available per month that they can use to ce corn, jeans, or a combination of both. The following table shows the amount of corn or jeans that can be produced using 1 hour of labor. Corn Jeans Dolorium 32 Arcadia 12 24 uses 1 million hours of labor per month to produce corm and 3 millon hours per month to produce jeans, while Dolonium uses per month to produce corn and 1 million hours per month to produce jeans. Consequently, Dolorium produces 24 million bushels of corn and 32 milion pairs of jeans, and Arcadia produces 12 millon bushels of corn and 72 million pairs of jeans. countries willing to trade goods, so, in the absence of trade between these two countries, each co produces. untry consumes the amount of corn and jeans it of jeans, and Arcadias opportunity cost of producing 1 bushel of corn is of jeans. Therefore, Y has a comparative advantage in the production of corn, and advantage in the production of jeans Suppose that each country completely specializes in the production of the good in which it has a comparative advantage, producing only that good. In this case, the country that produces corn will produce million bushels per month, and the country that produces jeans will produce million pairs per month. on the third row of the table (marked In the MacBook Pro
In the following table, enter each countrys production decision on the third row of the table (marked Production). Suppose the country that produces com trades 26 million bushels of corn to the other country in exchange for 78 million pairs of jeans. In the following table, select the amount of each good that each country exports and imports in the boxes across the row marked Trade Action, and enter each countrys final consumption of each good on the line marked Consumption. When the two countries did not specialize, the total production of com was 36 million bushels per month, and the total production of jeans was 104 million bushels per month, and the total production of yeans has increased by table. In the following table, enter this difference in the boxes Dolorium Jeans Corn Jeans s of pairs) ( Production 24 12 72 24 With Trade Production Trade action
Answer #1

Dolorium's opportunity cost of producing 1 bushel of corn is 32/8 = 4 pairs of jeans and Arcadia's opportunity cost of producing 1 bushel of corn is 24/12 =2 pairs of jeans. Because Arcadia has a lower opportunity cost for the production of corn. Therefore, Arcadia has a comparative advantage in the production of corn and Dolorium has a comparative advantage in the production of jeans.

Suppose that each country completely specializes in the production of the good in which it has a comparative advantage , producing only that good . In this case,the country that produces corn i.e Arcadia will produce (4)(12) = 48 million bushels per month And the country that produces jeans i.e Dolorium will produce 4(32)=128 million pairs per month.

Suppose the country that produces corn trades 26 million bushels of corn to the other country in exchange for 78 million pairs of jeans.

DOLORIUM ARCADIA
Corn (milions of bushels) Jeans (Millions of pairs) Corn (Millions of bushels) Jeans (Millions of pairs)
Without trade Production 24 32 12 72
Consumption 24 32 12 72
With Trade Production 0 128 48 0
Trade action Import 26 Export 78 Export 26 Import 78
Consumption 26 (128-78)=50 (48-26)=22 78
Gains from trade Increase in consumption (26-24)= 2 (50-32)= 18 (22-12)=10 (78-72)= 6

When the two countries did not specialize ,the total production of corn was 36 million bushels per month and the total production of jeans was 104 millions pairs per month. Because of specialization ,the total production of corn was 48 million bushels per month and total production of jeans was 128 million pair per month. This implies that because of specialization , total production of corn has increased by (48-36)=12 million bushels per month and total production of jeans has increased by (128-104)=24 million pairs per month.